If alpha and beta are the roots of the equation 2x^2+5x+3=0, find
(a)alpha^2+beta^2
(b) 2÷alpha+2÷beta
These question is written in words because alpha and beta signs aren't provided while typing.

Answer by rothauserc(4718)

2x^2 +5x +3 = 0
factor the polynomial
(2x+3) * (x+1) = 0
alpha = -3/2 and beta = -1
(a) (-3/2)^2 + (-1)^2 = (9/4) +(4/4) = 13/4
(b) (2/(-3/2)) + (2/-1) = (-1/3) + (-6/3) = -7/3